A ship travels 50 km on a bearing of 27 deg, then traveles on a bearing of 117 deg for 140 km. Find the distance traveled from the starting point to the ending point.
ashley917@netzero.net

Answer by stanbon(75887)

A ship travels 50 km on a bearing of 27 deg, then traveles on a bearing of 117 deg for 140 km. Find the distance traveled from the starting point to the ending point.
-----------------------------------------------------------
The x-coordinate of the final point is 50cos(27)+140cos(117)=-19
The y-coordinate of the final point is 50sin(27)+140sin(117)=23.67
distance from (0,0) to the final point is as follows:
d=sqrt((-19)^2+(23.6)^2)
=sqrt(921.27)
=30.35 km
------------------------------------------
Cheers,
Stan H.